OpenStudy (anonymous):

integral arctan sqrt x/(sqrt x(1+x))

OpenStudy (anonymous):

i know \[u=\arctan \sqrt{x} \]

OpenStudy (anonymous):

and \[du= 1/ \sqrt{x} (2x+2) dx\]

OpenStudy (helder_edwin):

do u mean \[\large \int\frac{\arctan\sqrt{x}}{\sqrt{x+x^2}}\,dx \] ?

OpenStudy (anonymous):

no, i mean \[\int\limits\limits \arctan \sqrt{x} / \sqrt{x} (1+x) dx\]

OpenStudy (anonymous):

\[\int\limits \frac{ \arctan \sqrt{x}}{ \sqrt{x}(1+x) } dx\]

OpenStudy (helder_edwin):

u were right \[\large u=\arctan\sqrt{x} \] so \[\large du=\frac{1}{1+x}\cdot\frac{1}{2\sqrt{x}}\,dx \] so \[\large 2du=\frac{dx}{\sqrt{x}(1+x)} \] so the integral becomes \[\large \int u\,2du=2\frac{u^2}{2}=u^2=\arctan^2\sqrt{x}+C \]

OpenStudy (anonymous):

thanks!

OpenStudy (helder_edwin):

u r welcome